OpenStudy (anonymous):

A dice is rolled. A card is drawn from a deck of face cards (i.e. only Jack, Queen, King of each suit). A coin is flipped. You win the game if you toss tails, draw a heart, and roll an even number. Find the probability that you do not win.

OpenStudy (anonymous):

First find the probability that you win. This is (3/12)*(1/2)*(3/6)=9/144=1/16 The probability that you do not win is 1-prob(win)=1-(1/16)=15/16 or a 94% chance you lose